OpenStudy (arhin):

show that In(1+x)=x-x^2/2+x^3/3-x^4/4+⋯

OpenStudy (tkhunny):

Can we assume \(\dfrac{1}{1-x} = 1 + x + x^{2} + x^{3} + ...\)?

OpenStudy (anonymous):

what @tkhunny said take the derivative, integrate term by term

OpenStudy (arhin):

ok

OpenStudy (tkhunny):

Well, you may wish to substute -x, first, giving \(\dfrac{1}{1+x} = 1 - x + x^{2} - x^{3} + ...\)